# Perverse microsheaves

### v2

@@ 1. Introduction
(1) $\dim_{\mathbb{C}}\{\,p\in M\,|\,H^{i}(p^{*}F)\neq 0\}\leq i\qquad\qquad\dim_{\mathbb{C}}\{\,p\in M\,|\,H^{i}(p^{!}F)\neq 0\}\leq i$
@@
Here $p^{*}$ and $p^{!}$ are the restriction functors to the point $p$.
@@
Perverse sheaves have played a pivotal role in many results in algebraic geometry and geometric representation theory. They turn to be natural both in terms of considerations of Frobenius eigenvalues in positive characteristic [3], and in terms of analytic considerations in characteristic zero, where they are the sheaves of solutions to regular holonomic differential equations, or more generally, D-modules [10, 21]. This latter equivalence highlights a key feature, not immediately apparent from the definition: despite being a seemingly arbitrarily demarcated subcategory of a category of complexes, perverse sheaves form an abelian category.
@@
There are microlocal versions (living on $T^{*}M$ or $\mathbb{P}T^{*}M$) of the category of $D$-modules [25, 12], perverse sheaves [2, 27], and the equivalence between them [1, 23, 28]. More generally still, Kashiwara has constructed a sheaf of categories on any complex contact manifold, locally equivalent to the microlocalization of $D$-modules [11], see also [24]. The variant appropriate to conic complex symplectic geometry has allowed the methods of geometric representation theory to be extended beyond cotangent bundles to more general symplectic resolutions and similar spaces [14, 4].
@@
The purpose of the present article is to construct categories of complex-constructible microsheaves on complex contact manifolds which naturally carry a perverse $t$-structure, globalizing the construction of Waschkies [27]. In the sequel we establish a Riemann-Hilbert equivalence with the canonical stack of $\mathcal{E}$-modules defined by Kashiwara [11].
@@
Our starting point is the globalization [26, 22] of the microlocal sheaf theory of Kashiwara and Schapira [15]. We recall the relevant notions in Section 4. In brief, the theory takes as input a real contact or exact symplectic manifold $V$, a choice of symmetric monoidal stable presentable coefficient category $\mathcal{C}$, and certain topological ‘Maslov data’. Maslov data is by definition a choice of trivialization for a certain canonical obstruction $V\to B^{2}Pic(\mathcal{C})$, and in fact agrees with the choices made in defining Floer-theoretic invariants in the same target spaces. The output of the theory is a sheaf of stable categories $\mu sh_{V}$ on $V$. For a locally closed subset $X\subset V$, we write $\mu sh_{X}\subset\mu sh_{V}|_{X}$ for the subsheaf of full subcategories on objects locally supported in $X$. A Legendrian or conic Lagrangian $L\subset V$ determines an obstruction $L\to BPic(\mathcal{C})$, a choice of trivialization for which (‘secondary Maslov data’) yields an equivalence $\mu sh_{L}\cong Loc_{L}$ with the sheaf of categories of local systems along $L$. Here we show that in the complex case:
Theorem 1.1 (see Section 3).
@@
Let $W$ be an exact complex symplectic manifold, and $\mathcal{C}$ the category of modules over a (discrete) commutative ring $R$. Then there is a canonical Maslov datum for $W$, with respect to which secondary Maslov data for a Lagrangian $L\subset W$ are identified with spin structures on $L$.
From this and the existing theory [22], we deduce:
Theorem 1.2 (see Section 6).
@@
Let $W$ be an exact complex symplectic manifold. For a discrete commutative ring $R$, there is a canonical sheaf of stable $R$-linear categories $\mu sh_{W}$ on $W$. For a complex Lagrangian $L\subset X$, a spin structure on $L$ determines an equivalence $\mu sh_{L}\cong Loc_{L}$.
Taking $D$ to be a disk around a smooth Lagrangian point $x\in X$, we obtain a canonical (up to noncanonical automorphism) ‘microstalk’ functor $\mu_{x}:\mu sh_{X}(W)\to\mu sh_{X}(D)\xrightarrow{\sim}R-mod$.
@@
We write $\mu sh_{W,\mathbb{C}-c}\subset\mu sh_{W}$ for the sheaf of full subcategories on objects whose (micro)support is a complex analytic Lagrangian subset of $W$. We define $(\mu sh_{W,\mathbb{C}-c})^{\leq 0}\subset\mu sh_{W,\mathbb{C}-c}$ as the sheaves of full subcategories on those objects all of whose microstalks, as elements of $R-mod$, have cohomology concentrated in degrees $\leq 0$. We define similarly $(\mu sh_{W,\mathbb{C}-c})^{\geq 0}$.
@@
To introduce these other categories, we return to the basic definitions of microsheaves. For $M$ a manifold, we have the presheaf of categories $\mu sh^{pre}$ on $T^{*}M$ defined by
$\mu sh^{pre}(U):=sh(M)/\{F\,|\,ss(F)\cap U=\emptyset\}.$
We write $\mu sh$ for its sheafification.
Let us write $T^{*}M_{\mathbb{R}}$ for the topological space whose open sets are the $\mathbb{R}^{>0}$-invariant open subsets of $T^{*}M$, and $\gamma_{\mathbb{R}}:T^{*}M\to T^{*}M_{\mathbb{R}}$ for the continous map given by the set theoretic identity. Let us write $\mu_{\mathbb{R}}sh$ for the sheafification of $(\gamma_{\mathbb{R}})_{*}\mu sh^{pre}$. Because microsupports are conic, there is an isomorphism of the sheaf pullback $\gamma_{\mathbb{R}}^{*}\mu_{\mathbb{R}}sh\xrightarrow{\sim}\mu sh$. In fact, because $\mathbb{R}^{>0}$ is contractible, the presheaf preimage $\gamma_{\mathbb{R}}^{*,pre}\mu_{\mathbb{R}}sh$ is already a sheaf, so already isomorphic to $\mu sh$. That is, nothing is gained or lost by considering $\mu_{\mathbb{R}}sh$ instead of $\mu sh$.
By contrast, when $M$ is a complex manifold, consider similarly the topological space $T^{*}M_{[\mathbb{C}^{*}]}$ on $\mathbb{C}^{*}$-invariant open sets in $T^{*}M$, and define $\mu_{\mathbb{C}}sh$ etc. similarly. Now there are maps
$\gamma_{\mathbb{C}}^{*,pre}\mu_{\mathbb{C}}sh\to\gamma_{\mathbb{C}}^{*}\mu_{\mathbb{C}}sh\to\mu sh$
These maps are all fully faithful on sections [27, Sec. 2.4], but we do not know when or whether (or after imposing what microsupport conditions) they are essentially surjective.
Consider the projectivized cotangent bundle, $\pi:T^{*}M\setminus 0_{M}\to\mathbb{P}^{*}M$. Then we have an inclusion of sheaves of full subcategories:
(2) $\mathbb{P}\mu sh:=\pi_{*}\gamma_{\mathbb{C}^{*}}^{*,pre}\mu_{\mathbb{C}}sh\hookrightarrow\pi_{*}\mu sh$
More generally, for $V$ a complex contact manifold and $\pi:\tilde{V}\to V$ its symplectization, we will identify a sheaf of full subcategories $\mathbb{P}\mu sh_{V}\subset\pi_{*}\mu sh_{\tilde{V}}$ which agrees with (2) in local charts (see Definition 5.14). We write $\mathbb{P}\mu sh_{V,\mathbb{C}-c}\subset\mathbb{P}\mu sh_{V}$ for the subsheaf of full subcategories on objects with complex Legendrian microsupport; note $\mathbb{P}\mu sh_{V,\mathbb{C}-c}=\mathbb{P}\mu sh_{V}\cap\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c}$.
For complex exact symplectic manifolds $(W,\lambda)$, we may use the contactization $(W\times\mathbb{C},\lambda+dz)$ to define a sheaf of categories $\mathbb{P}\mu sh_{W\times 0}$. We do not understand the relationship of this with $\mu sh_{W}$ in general. However, if the Liouville flow on $(W,\lambda)$ integrates to a weight-1 $\mathbb{C}^{*}$-action, then $\mathbb{C}^{*}$ naturally acts on $W\times\mathbb{C}$ by contactomorphism. Let $\gamma_{\mathbb{C}}$ be the set-theoretic identity on $W\times\mathbb{C}^{*}$ (resp. on $W$) where the source carries the Euclidean topology but the target is endowed with the $\mathbb{C}^{*}$ invariant topology. In Theorem 6.3, we show that there is a natural $\mathbb{Z}[t,t^{-1}]$-linear structure on $\lambda_{*}\mathbb{P}\mu sh_{W\times\{0\}}$, and there are fully faithful embeddings
(3) $(\gamma_{\mathbb{C}})_{*}\mathbb{P}\mu sh_{W\times\{0\}}\hookrightarrow(\gamma_{\mathbb{C}})_{*}\mu sh_{W}\otimes\mathbb{Z}[t,t^{-1}]$
(4) $\mu_{\mathbb{C}}sh_{W}:=(\gamma_{\mathbb{C}})_{*}\mathbb{P}\mu sh_{W\times\{0\}}\otimes_{\mathbb{Z}[t,t^{-1}]}\mathbb{Z}\hookrightarrow(\gamma_{\mathbb{C}})_{*}\mu sh_{W}$
We can now state our results on $t$-structures.
@@
Let $V$ be a complex contact manifold. Then the pair
$\mathbb{P}\mu sh_{V,\mathbb{C}-c}^{\geq 0}:=\mathbb{P}\mu sh_{V}\cap\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c}^{\geq 0}\qquad\qquad\qquad\mathbb{P}\mu sh_{V,\mathbb{C}-c}^{\leq 0}:=\mathbb{P}\mu sh_{V}\cap\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c}^{\leq 0}$
determines a $t$-structure on $\mathbb{P}\mu sh_{V,\mathbb{C}-c}$.
@@
Let $W$ be a complex exact symplectic manifold whose Liouville vector field integrates to a weight-1 $\mathbb{C}^{*}$ action. Then the pair
$\mu_{\mathbb{C}}sh_{W,\mathbb{C}-c}^{\geq 0}:=\mu_{\mathbb{C}}sh_{W}\cap\mu sh_{W,\mathbb{C}-c}^{\geq 0}\qquad\qquad\qquad\mu_{\mathbb{C}}sh_{W,\mathbb{C}-c}^{\leq 0}:=\mu_{\mathbb{C}}sh_{W}\cap\mu sh_{W,\mathbb{C}-c}^{\leq 0}$
determines a $t$-structure on $\mu_{\mathbb{C}}sh_{W,\mathbb{C}-c}$. Moreover, the Hom sheaf of two objects in the heart is a ($\frac{1}{2}\dim W$-shifted) perverse sheaf.
Example 1.5.
Take $(W,\lambda)$ as in Theorem 1.4, and suppose in addition $(W,\operatorname{re}\lambda)$ is a Weinstein manifold admiting a complete collection of generalized cocores, so that the main theorem of [7] can be applied. Many examples coming from geometric representation theory – quiver varieties, hypertoric varieties, moduli of Higgs bundles, etc. satisfy this assumption; cf. [29]. Let $L$ be the core of $W$. Then $\mu sh_{L}(W)$ is equivalent to (the module category of) $Fuk(W)$ [7], and so Theorem 1.4 asserts we have constructed a $t$-structure on the full subcategory
$\mu_{\mathbb{C}}sh_{L}(W)\hookrightarrow\mu sh_{L}(W)\cong Fuk(W)$
and correspondingly the perverse microsheaves determine a full abelian subcategory of $H^{0}Fuk(W)$.
Let us conclude by explicitly advertising the question:
Definition 1.6.
Let $W$ be a complex exact symplectic manifold whose Liouville vector field integrates to a weight-1 $\mathbb{C}^{*}$ action. We say a complex conic Lagrangian $L\subset W$ has only complex global microsheaves if $\mu_{\mathbb{C}}sh_{L}(W)\hookrightarrow\mu sh_{L}(W)$ is essentially surjective.
Question 1.7.
Which conic complex Lagrangians have only complex global microsheaves?
We do not know that the answer is not “all of them”. (In fact, in a previous draft of this article, we blithely assumed this was the case.) Example 1.5 provides one motivation for the question, since in the notation there, whenever $L$ has only complex global microsheaves, we have constructed a $t$-structure on the Fukaya category of $W$.
Here is what little we know about 1.7. It is not difficult to see that $\mathbb{C}^{*}$-conic subsets of cotangent bundles are completely complex, and slightly more generally, that the same holds for Lagrangians which admit a cover by $\mathbb{C}^{*}$-conic subsets of cotangent bundles. (This implies the existence of $t$-structures on Fukaya categories of hypertoric varieties, ALE spaces, and a few other quiver varieties.) We do not presently know much more in general.
Acknowledgements. We wish to thank Roman Bezrukavnikov, Sanath Devalapurkar, Peter Haine, Benjamin Gammage, Justin Hilburn, Michael McBreen, Semon Rezchikov, Marco Robalo, Pierre Schapira and Filip Živanović for enlightening conversations.
LC was partially supported by Simons Foundation grant 385573 (Simons Collaboration on Homological Mirror Symmetry). DN was partially supported by NSF grant DMS-2101466. CK and VS were partially supported by: VILLUM FONDEN grant VILLUM Investigator 37814, and VS was supported in addition by Novo Nordisk Foundation grant NNF20OC0066298, Danish National Research Foundation grant DNRF157, and the USA NSF grant CAREER DMS-1654545.
@@ 2. Complex contact and symplectic manifolds
A complex (or holomorphic) symplectic manifold is a complex manifold $X$ along with a closed holomorphic $2$-form $\omega\in H^{0}(X,\Omega^{2}_{X})$.
A complex symplectic manifold $(X,\omega)$ determines a family, parameterized by $\hbar\in\mathbb{C}^{*}$, of real symplectic manifolds $(X,\operatorname{re}(\hbar\omega))$. By an exact complex symplectic manifold, we mean a pair $(X,\lambda)$ where $X$ is a complex manifold and $\lambda$ is a holomorphic 1-form such that $d\lambda=\partial\lambda$ is symplectic.
@@
Let $X$ be a complex manifold. Then the holomorphic cotangent bundle $\Omega_{X}$ carries a canonical holomorphic $1$-form $\lambda_{can,\mathbb{C}}=ydx$, where $(x_{1},\dots,x_{n},y_{1},\dots,y_{n})$ are canonical holomorphic coordinates. Then $(\Omega_{X},d\lambda_{can,\mathbb{C}})$ is a complex symplectic manifold.
Meanwhile, the real cotangent bundle $T^{*}M$ carries the canonical $1$-form $\lambda_{can}$. There is a natural identification $\Omega_{X}\to T^{*}M$ defined in local holomorphic coordinates by
@@
A complex (or holomorphic) contact manifold is a complex manifold $V$ of complex dimension $12n+$ 1 along with a holomorphic hyperplane field $\mathcal{H}\hookrightarrow\mathcal{T}_{V}$ which is maximally non-integrable. Concretely, this means that if $\alpha\in\Omega_{V}(U)$ is a holomorphic $1$-form for which $\mathcal{H}=\ker\alpha$ in some local chart $U\subset V$, then $\alpha\wedge(d\alpha)^{n}\neq 0$.
Given a complex contact manifold $(V,\xi)$, there is a holomorphic line bundle $\mathcal{T}_{V}/\mathcal{H}\to V$. Contact forms are naturally sections of $(\mathcal{T}_{V}/\mathcal{H})^{\vee}$. Note the existence of holomorphic global sections is a rather more stringent condition than the corresponding condition in real smooth geometry; correspondingly complex contact manifolds typically do not have global contact forms.
@@
$\pi:\tilde{V}:=(\mathcal{T}_{V}/\mathcal{H})^{\vee}\setminus 0_{V}\to V$
which we call the complex symplectization of $V$ and let $\lambda_{\tilde{V}}$ denote the pullback of $\lambda_{can,\mathbb{C}}$ under the inclusion $\tilde{V}\hookrightarrow\Omega_{V}$.
We also consider the projectivized $S^{1}$-bundle $p:\tilde{V}/\mathbb{R}_{+}\to V$ and set $\xi_{\hbar}:=\ker(\operatorname{re}(\hbar\lambda_{\tilde{V}}))$ for $\hbar\in\mathbb{C}^{*}$. The relation between these spaces is summarized by the following diagram:
@@
    $(\tilde{V},\hbar\lambda_{\tilde{V}})$ is an exact complex symplectic manifold
@@
    $(\tilde{V}/\mathbb{R}_{+},\xi_{\hbar})$ is a real contact manifold
@@
    $(\tilde{V},\operatorname{re}(\hbar\lambda_{\tilde{V}}))$ is canonically isomorphic to the real symplectization of $(\tilde{V}/\mathbb{R}_{+},\xi_{\hbar})$. (This isomorphism intertwines the $1$-forms and the $\mathbb{R}_{+}$-bundle structure over $\tilde{V}/\mathbb{R}_{+}$).
@@
We compute $\hbar\lambda_{\tilde{V}}$ locally by choosing a holomorphic contact $1$-form $\alpha$ defined on some open set $U\subset V$. Such a choice induces a holomorphic embedding
$\iota_{\alpha}:\mathbb{C}^{*}\times U$ $\hookrightarrow\tilde{V}\subset\Omega_{V}$
@@
To compute the pullback of $\hbar\lambda_{\tilde{V}}$ under $\iota_{\alpha}$, choose $(z,x)\in\mathbb{C}^{*}\times U$ and $d\iota_{\alpha}(v)\in T\tilde{V}_{z\alpha_{x}}\subset T\Omega_{V}$. Then we have $\hbar\lambda_{\tilde{V}}(d\iota_{\alpha}(v))=\hbar\lambda_{can,\mathbb{C}}(d\iota_{\alpha}(v))=z\hbar\alpha_{x}(d\pi\circ d\iota_{\alpha}(v))$. Hence
(6) $\iota_{\alpha}^{*}\hbar\lambda_{\tilde{V}}=z\hbar\alpha.$
Both (i) and (ii) can be checked immediately using (6). The proof of (iii) is an exercise in chasing definitions so is omitted.
@@
The complex projectivization $V:=(\Omega_{X}-0_{X})/\mathbb{C}^{*}$ is a complex contact manifold with respect to $\lambda_{can,\mathbb{C}}$ and we have $\tilde{V}=(\Omega_{X}-0_{X})$. The associated real projectivization $\tilde{V}/\mathbb{R}_{+}$ carries a circle of real contact forms $\operatorname{re}(\hbar\lambda_{can,\mathbb{C}})$.

### v3

@@ 1. Introduction
(1) $\dim_{\mathbb{C}}\{\,x\in M\,|\,H^{i}(\iota_{x}^{*}F)\neq 0\}\leq-i\qquad\qquad\dim_{\mathbb{C}}\{\,x\in M\,|\,H^{i}(\iota_{x}^{!}F)\neq 0\}\leq i$
@@
Here we denote by $\iota_{x}:\{x\}\hookrightarrow M$ the inclusion of the point $p$ to $M$ and by $\iota_{x}^{*}$ and $\iota_{x}^{!}$ the induced restriction functors.
@@
Perverse sheaves have played a pivotal role in many results in algebraic geometry and geometric representation theory. They turn out to be natural both in terms of considerations of Frobenius eigenvalues in positive characteristic [3], and in terms of analytic considerations in characteristic zero, where they are the sheaves of solutions to regular holonomic differential equations, or more generally, D-modules [19, 23, 20, 34]. This latter equivalence highlights a key feature, not immediately apparent from the definition: despite being a seemingly arbitrarily demarcated subcategory of a category of complexes, perverse sheaves form an abelian category.
@@
There are microlocal versions (living on $T^{*}M$ or $\mathbb{P}T^{*}M$) of the category of $D$-modules [38, 22], perverse sheaves [2, 42], and the equivalence between them [1, 36, 43]. More generally still, Kashiwara has constructed a sheaf of categories on any complex contact manifold, locally equivalent to the microlocalization of $D$-modules [21], see also [37]. A variant of this construction appropriate to conic complex symplectic geometry has allowed the methods of geometric representation theory to be extended beyond cotangent bundles to more general symplectic resolutions and similar spaces [24, 4].
@@
The purpose of the present article is to construct perverse $t$-structures on categories of complex-constructible microsheaves, globalizing the construction of Waschkies [42]. In the sequel [29] we establish a Riemann-Hilbert equivalence with the canonical stack of $\mathcal{E}$-modules defined by Kashiwara [21].
@@
Our starting point is the globalization [39, 35] of the microlocal sheaf theory of Kashiwara and Schapira [25]. We recall the relevant notions in Section 4. In brief, the theory takes as input a real contact or exact symplectic manifold $V$, a choice of symmetric monoidal stable compactly generated coefficient category $\mathcal{C}$, and a trivialization of a certain canonical obstruction $V\to B^{2}Pic(\mathcal{C})$. We refer to said trivialization as a Maslov datum; is is also what is required to define Floer-theoretic invariants in the same target spaces. The output of the theory is a sheaf of stable categories $\mu sh_{V}$ on $V$ [35, Thm. 1.1]. For a locally closed subset $X\subset V$, we write $\mu sh_{X}\subset\mu sh_{V}|_{X}$ for the subsheaf of full subcategories on objects locally supported in $X$.
A Legendrian or conic Lagrangian $L\subset V$ determines an obstruction $L\to BPic(\mathcal{C})$, a choice of trivialization for which (a ‘secondary Maslov datum’) yields an equivalence
$\mu sh_{L}\cong loc_{L}$
with the sheaf of categories of local systems along $L$ [35, Thm. 1.2]. In particular, if $D\subset V$ is a smooth Legendrian disk containing a point $p$, a choice of secondary Maslov datum for $D$ determines an equivalence
(2) $(\mu sh_{D})_{p}\xrightarrow{\sim}\mathcal{C}.$
The space of secondary Maslov data for the disk $D$ is a torsor for $BPic(\mathcal{C})$, which acts on maps (2) in the evident way.
In this paper, we will be interested in contact and conic symplectic manifolds which come from complex geometry. It turns out that such manifolds admit a canonical choice of Maslov datum, yielding a canonical notion of $\mu sh$. For simplicity, we state this in the symplectic setting. By an exact complex symplectic manifold, we mean a complex manifold $W$ along with a holomorphic $1$-form $\lambda$ such that $d\lambda$ is holomorphic symplectic. The underlying real manifold of $W$ carries the real exact symplectic structure $\operatorname{re}(\lambda)$, so that we can meaningfully discuss Maslov data and microlocal sheaves on $W$.
Theorem 1.1.
@@
Let $W$ be an exact complex symplectic manifold and $\mathcal{C}$ the category of modules over a (discrete) commutative ring $R$. Then there is a canonical Maslov datum for $W$, with respect to which secondary Maslov data for a conic complex Lagrangian $L\subset W$ are identified with $R$-spin structures on $L$.
By an $R$-spin structure on a real symplectic manifold $W$, we mean a null-homotopy of the composition $W\to BU\xrightarrow{w_{2}}B^{2}\mathbb{Z}^{\times}\to B^{2}R^{\times}$. When $R=\mathbb{Z}$, this is a spin structure in the ordinary sense; in general, if such structures exist, they form a torsor for $\mathrm{H}^{1}(X,R^{\times})$. The proof of Theorem 1.1 is in Section 4.6, where we also deduce, from the aforementioned general properties of microlocal sheaves:
Corollary 1.2 (The canonical microsheaf category).
@@
Let $W,\mathcal{C}$ be as in Theorem 1.1. There is a canonical sheaf of stable $\mathcal{C}$-linear categories $\mu sh_{W}$ on $W$. For a complex conic Lagrangian $L\subset W$, an $R$-spin structure $\sigma$ on $L$ determines an equivalence $\mu sh_{L}\cong loc_{L}$.
In particular, specializing to the neighborhood of a point gives:
Corollary 1.3 (The microstalk functor).
Let $W,\mathcal{C}$ be as in Theorem 1.1. If $X\subseteq W$ is a closed subset which locally around $p\in X$ is conic complex Lagrangian, there is a functor
(3) $\omega_{p}^{-1}:(\mu sh_{X})_{p}\xrightarrow{\sim}\mathcal{C},$
which is well-defined up to non-canonical invertible natural transformation. In particular, given $K\in\mu sh_{X}(W)$, the isomorphism class of $\omega_{p}^{-1}(K)$ is well defined.
We refer to $\omega_{p}^{-1}(K)$ as the microstalk of $K$ at $p$. In Section 5.2, we will give a more explicit characterization of this microstalk.
Definition 1.4.
@@
Write $\mu sh_{W,\mathbb{C}-c}\subset\mu sh_{W}$ for the sheaf of full subcategories on objects whose (micro)support is a complex analytic Lagrangian subset of $W$. We define
(4) $(\mu sh_{W,\mathbb{C}-c})^{\leq 0}\subset\mu sh_{W,\mathbb{C}-c}\qquad\qquad(\mu sh_{W,\mathbb{C}-c})^{\geq 0}\subset\mu sh_{W,\mathbb{C}-c}$
as the sheaves of full subcategories on those objects all of whose microstalks, as elements of RmodR-mod, have cohomology concentrated in degrees $\leq 0$, resp. $\geq 0$.
@@
Recall that for a complex contact manifold $V$, its symplectization is the $\mathbb{C}^{*}$ bundle $\pi:\widetilde{V}\to V$ whose (local) holomorphic sections give (local) complex contact forms; it carries canonically an exact complex symplectic form (as we review in more detail in Section 2) and hence we have the canonically defined $\mu sh_{\widetilde{V}}$.
@@
Then the pair $\left((\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c})^{\leq 0},(\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c})^{\geq 0}\right)$ determines a $t$-structure on $\pi_{*}\mu sh_{\widetilde{V},\mathbb{C}-c}$ (i.e., the sections of the above sheaves over any open set $U$ give a t-structure on $\pi_{*}\mu sh_{\tilde{V},\mathbb{C}-c}(U)$).
For an exact complex symplectic manifold $(W,\lambda)$, we may use the contactization $(W\times\mathbb{C},\lambda+dz)$ to define a sheaf of categories $\pi_{*}\mu sh_{\pi^{-1}(W\times\{0\})}$, i.e. the sheaf of full subcategories of $\pi_{*}\mu sh_{\widetilde{W\times\mathbb{C}}}$ on objects whose support is contained in $\pi^{-1}(W\times\{0\})$. We do not understand the relationship of this with $\mu sh_{W}$ in general. However, if the Liouville flow on $(W,\lambda)$ integrates to a weight-1 $\mathbb{C}^{*}$-action, then $\mathbb{C}^{*}$ naturally acts on $W\times\mathbb{C}$ by contactomorphism. Let $\gamma_{\mathbb{C}}$ be the set-theoretic identity on $W\times\mathbb{C}$ (resp. on $W$) where the source carries the Euclidean topology but the target is endowed with the $\mathbb{C}^{*}$ invariant topology.
@@
Let $W$ be a complex exact symplectic manifold whose Liouville vector field integrates to a weight-1 $\mathbb{C}^{*}$ action. Then the pair $((\gamma_{\mathbb{C}})_{*}\mu sh_{W,\mathbb{C}-c})^{\geq 0},((\gamma_{\mathbb{C}})_{*}\mu sh_{W,\mathbb{C}-c})^{\leq 0}$ determines a $t$-structure on $(\gamma_{\mathbb{C}})_{*}\mu sh_{W,\mathbb{C}-c}$. Moreover, the Hom sheaf of two objects in the heart is a ($\frac{1}{2}\dim W$-shifted) perverse sheaf.
Finally, we recall that there is a comparison theorem [11] between microsheaves and Fukaya categories; consequently, our results can be translated to give $t$-structures on certain Fukaya categories. We spell this out in Appendix B.
Acknowledgements. We wish to thank Roman Bezrukavnikov, Sanath Devalapurkar, Peter Haine, Benjamin Gammage, Sam Gunningham, Justin Hilburn, Wenyuan Li, Michael McBreen, Semon Rezchikov, Marco Robalo, Pierre Schapira and Filip Živanović for enlightening conversations. We are particularly grateful to Sanath Devalapurkar for help with Section 3 and to Wenyuan Li for help with Proposition 5.9.
LC was partially supported by Simons Foundation grant 385573 (Simons Collaboration on Homological Mirror Symmetry). DN was partially supported by NSF grant DMS-2101466. CK and VS were partially supported by: VILLUM FONDEN grant VILLUM Investigator 37814, and VS was supported in addition by Novo Nordisk Foundation grant NNF20OC0066298, Danish National Research Foundation grant DNRF157, and the USA NSF grant CAREER DMS-1654545. CK was supported by Max Planck Institude for Mathematics in Bonn during the revision of the paper.
@@ 2. Complex contact and symplectic manifolds
A complex (or holomorphic) symplectic manifold is a complex manifold $W$ along with a closed holomorphic $2$-form $\omega\in H^{0}(W,\Omega^{2}_{W})$.
A complex symplectic manifold $(W,\omega)$ determines a family, parameterized by $\hbar\in\mathbb{C}^{*}$, of real symplectic manifolds $(W,\operatorname{re}(\hbar\omega))$. By an exact complex symplectic manifold, we mean a pair $(W,\lambda)$ where $W$ is a complex manifold and $\lambda$ is a holomorphic 1-form such that $d\lambda=\partial\lambda$ is symplectic.
@@
Let $M$ be a complex manifold. Then the holomorphic cotangent bundle $\Omega_{M}$ carries a canonical holomorphic $1$-form $\lambda_{can,\mathbb{C}}=ydx$, where $(x_{1},\dots,x_{n},y_{1},\dots,y_{n})$ are canonical holomorphic coordinates. Then $(\Omega_{M},d\lambda_{can,\mathbb{C}})$ is a complex symplectic manifold.
Meanwhile, the real cotangent bundle $T^{*}M$ carries the canonical $1$-form $\lambda_{can}$. There is a natural identification $\Omega_{M}\to T^{*}M$ defined in local holomorphic coordinates by
@@
A complex (or holomorphic) contact manifold is a complex manifold $V$ of complex dimension $2n+1$ along with a holomorphic hyperplane field $\mathcal{H}\hookrightarrow\mathcal{T}_{V}$ which is maximally non-integrable. Concretely, this means that if $\alpha\in\Omega_{V}(U)$ is a holomorphic $1$-form for which $\mathcal{H}=\ker\alpha$ in some local chart $U\subset V$, then $\alpha\wedge(d\alpha)^{n}\neq 0$.
Given a complex contact manifold $(V,\xi)$, there is a holomorphic line bundle $\mathcal{T}_{V}/\mathcal{H}\to V$. (Local) complex contact forms are nonvanishing (local) nonvanishing holomorphic sections of $(\mathcal{T}_{V}/\mathcal{H})^{\vee}$. Correspondingly, a global complex contact form is a global holomorphic trivialization of this line bundle (which does not typically exist).
@@
$\pi:\widetilde{V}:=(\mathcal{T}_{V}/\mathcal{H})^{\vee}\setminus 0_{V}\to V$
which we call the complex symplectization of $V$ and let $\lambda_{\widetilde{V}}$ denote the pullback of $\lambda_{can,\mathbb{C}}$ under the inclusion $\widetilde{V}\hookrightarrow\Omega_{V}$.
We also consider the projectivized $S^{1}$-bundle $p:\widetilde{V}/\mathbb{R}_{+}\to V$ and set $)\xi_{\hbar}:=\ker(\operatorname{re}(\hbar\lambda_{\widetilde{V}}$)) for $\hbar\in\mathbb{C}^{*}$. The relation between these spaces is summarized by the following diagram:
@@
[…]
@@
    $(\widetilde{V}/\mathbb{R}_{+},\xi_{\hbar})$ is a real contact manifold
@@
    $)(\widetilde{V},\operatorname{re}(\hbar\lambda_{\widetilde{V}}$)) is canonically isomorphic to the real symplectization of $(\widetilde{V}/\mathbb{R}_{+},\xi_{\hbar})$. (This isomorphism intertwines the $1$-forms and the $\mathbb{R}_{+}$-bundle structure over $\widetilde{V}$$\mathbb{R}_{+}$).
@@
We compute $\hbar\lambda_{\widetilde{V}}$ locally by choosing a holomorphic contact $1$-form $\alpha$ defined on some open set $U\subset V$. Such a choice induces a holomorphic embedding
$\iota_{\alpha}:\mathbb{C}^{*}\times U$ $\hookrightarrow\widetilde{V}\subset\Omega_{V}$
@@
To compute the pullback of $\hbar\lambda_{\widetilde{V}}$ under $\iota_{\alpha}$, choose $(z,x)\in\mathbb{C}^{*}\times U$ and $d\iota_{\alpha}(v)\in T\widetilde{V}_{z\alpha_{x}}\subset T\Omega_{V}$. Then we have $\hbar\lambda_{\widetilde{V}}(d\iota_{\alpha}(v))=\hbar\lambda_{can,\mathbb{C}}(d\iota_{\alpha}(v))=z\hbar\alpha_{x}(d\pi\circ d\iota_{\alpha}(v))$. Hence
(6) $\iota_{\alpha}^{*}\hbar\lambda_{\widetilde{V}}=z\hbar\alpha.$
Both (i) and (ii) can be checked immediately using (6). The proof of (iii) is an exercise in chasing definitions. ∎
@@
The complex projectivization $V:=(\Omega_{X}-0_{X})/\mathbb{C}^{*}$ is a complex contact manifold with respect to $\lambda_{can,\mathbb{C}}$ and we have $\widetilde{V}=(\Omega_{X}-0_{X})$. The associated real projectivization $\widetilde{V}$$\mathbb{R}_{+}$ carries a circle of real contact forms $\operatorname{re}(\hbar\lambda_{can,\mathbb{C}})$.